OpenStudy (anonymous):

How to factor 4X^2 + 16X +15? Please show the steps!

OpenStudy (anonymous):

(2x+5)(2x+3) First you find two terms that you multiply to get the first 4x^2...you just have to try different things out sometimes until one works. And then you look at the last term of "+15" and see what two numbers you multiply to get that

OpenStudy (anonymous):

So you don't look at the middle term (16x)?

OpenStudy (anonymous):

You do but it's easier to find the terms to multiply since there are less numbers that multiply to 15 than numbers that add up to 16x if that makes sense

OpenStudy (anonymous):

I just list out numbers that could multiply to the last value "15" and try them out with the first variable "2x" in the factor to see if it'll add up to 16x

OpenStudy (anonymous):

4 × 15 = 60 60 = 10 × 6 10 + 6 = 16.

OpenStudy (anonymous):

^ is more reliable than 'guess-and-check.'

OpenStudy (anonymous):

4X^2 + 16X +15 = 4x^2 +10x +6x +15 = 2x(2x+5) + 3(2x+5) = (2x+5)(2x+3) No guessing required.

OpenStudy (anonymous):

@1223 Make sense?

OpenStudy (anonymous):

Yes it does! Thank You!